OpenStudy (anonymous):

Which exponential function goes through the points (1, 16) and (4, 128)? f(x) = 4(4)x f(x) = 8(2)x f(x) = 8(2)-x f(x) = 4(4)-x

OpenStudy (anonymous):

I think its B am I right?

OpenStudy (anonymous):

@darkside3704 @quetejedi @QueenBee232 @Rina.r @ash2326

OpenStudy (anonymous):

@cutecupcake

OpenStudy (dumbcow):

you are correct :)

OpenStudy (anonymous):

could you show me how to solve it thanks my brother gave me the answer but doesn't have the patience to teach me

OpenStudy (dumbcow):

haha yeah sure you start with general exponential function \[y = a(b^x)\] plug in the given points for x,y \[a*b = 16\] \[a*b^4 = 128\] solve for a,b by substitution from 1st equation, solve for a \[a = \frac{16}{b}\] sub into 2nd equation \[\frac{16}{b}*b^4 = 128\] solve for b \[16b^3 = 128\] \[b^3 = 8\] \[b = 2\] plug back in to get a \[a = \frac{16}{2} = 8\] \[y = 8(2^x)\]

OpenStudy (anonymous):

woaah :O

OpenStudy (anonymous):

thankyou can you help me solve another one?

OpenStudy (dumbcow):

ok

OpenStudy (anonymous):

The table below shows the values of f(n) for different values of n. n 1 2 3 4 5 6 f(n) 1 2 5 12 29 70 Which recursive function best represents the values shown in the table? f(1) = 1, f(2) = 2, f(n) = 2f(n -1) f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = f(n -3) + f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = 2f(n -1) + f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = f(n -3) f(n - 2); n > 2

OpenStudy (anonymous):

I think its c

OpenStudy (anonymous):

i agree

OpenStudy (dumbcow):

yep its c f(n-3) is impossible since f(0) is not defined

OpenStudy (anonymous):

chief ur doing good :)

OpenStudy (anonymous):

he function below shows the relationship between the length of each side of a square (y) and the area of the square (x+2).  Which graph best shows the function?